To purchase a new truck, you borrow $30,000. The bank offers a 6-year loan at an interest rate of 3.25% compounded annually. If you make only one payment at the end of the loan period, repaying the principal and interest:

a. What is the number of time periods (n) you should use in solving this problem?

b. What rate of interest (i), per period of time, should be used in solving this problem?

c. Is the present single amount of money (P) known? (Yes or No)

d. Which time value factor should be used to solve this problem?

e. What is the total amount that must be paid back?

f. How much of the total amount repaid represents interest?
Answer #1

Loan amount = $30,000. Time period = 6-year. Rate of interest = 3.25% compounded annually.

a. Number of time periods (n) = 6 years

b. Rate of interest (i), per period of time, = 3.25%

c.Yes it is $30000

d. Use (F/P, i%, n)

e Total amount that must be paid back = $30000 (F/P, 3.25%, 6) = 1.21155*30000 = $36,346.50

f. Interest amount = $36,346.50 - $30,000 = $6,346.50